OpenStudy (anonymous):

simplify : (√125y^6)^(1/3)

OpenStudy (saifoo.khan):

where does the root ends?

OpenStudy (anonymous):

oh sorry. It ends at 6. the square root is cube rooted.

OpenStudy (saifoo.khan):

\[\sqrt[3]{125y^6} \]

OpenStudy (saifoo.khan):

like that?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

NONONON

OpenStudy (saifoo.khan):

5y

OpenStudy (anonymous):

Its the cube root of the square root of 125y6

OpenStudy (saifoo.khan):

yes.

OpenStudy (saifoo.khan):

the answer is that^

OpenStudy (anonymous):

I think he means \[\sqrt[3]{\sqrt{125y^6}?}\]

OpenStudy (anonymous):

the answer says y√5 but i want to know how to do it..

OpenStudy (anonymous):

YESSS thats what i mean^^

OpenStudy (anonymous):

well the y is simple y\[y^3*y^3 \] you can pull one out and then pull a y out

OpenStudy (anonymous):

oh i see!

OpenStudy (saifoo.khan):

this?

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{125y^6}\]

OpenStudy (anonymous):

theres a square inside the cube root

OpenStudy (saifoo.khan):

\[\huge \sqrt[3]{\sqrt{125y^5}}\]

OpenStudy (anonymous):

yes but y^6

OpenStudy (saifoo.khan):

alright.

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

\[\Large \sqrt{5}y\]

OpenStudy (saifoo.khan):

tht's the answer.

OpenStudy (anonymous):

howd you do it?

OpenStudy (saifoo.khan):

the square root can be written as exponent 1/2 the cube root can be written as exponent 1/2

OpenStudy (anonymous):

cube root is 1/3